OpenStudy (anonymous):

(a) Let F = 5y i  + 2 j  −  k and C be the line from (3, 2, -2) to (6, 1, 7). Find integral_C F · dr= (b) Let G = 5xe^(5x^2+2y^2+z^2) i  + 2ye^(5x^2+2y^2+z^2) j  + ze^(5x^2+2y^2+z^2) k  and C be the same as in part (a). find integral_C G · dr =

OpenStudy (anonymous):

@ganeshie8 can you help me with this one for part a i use the F.T. F.L.I and i get -5 but its wrong

OpenStudy (anonymous):

@ganeshie8 I did it again and i got -11 but its also wrong. what am I doing wrong

OpenStudy (anonymous):

gradf(5yi+2j-k) = 5yx+2y -z right??

ganeshie8 (ganeshie8):

you need to parameterize and setup the line integral, right ?

OpenStudy (anonymous):

i'm not sure

OpenStudy (anonymous):

for part a i try using the FTFLT

ganeshie8 (ganeshie8):

for part b, try \(\large \frac{e^{231}-e^{57}}{2}\)

OpenStudy (anonymous):

its right, how did you figured it out?

ganeshie8 (ganeshie8):

vector field in part b is conservative so you can eyeball the potential function

ganeshie8 (ganeshie8):

find the curl

ganeshie8 (ganeshie8):

you should get 0 if the field is conservative

OpenStudy (anonymous):

what do you do after you find that the curl is 0

ganeshie8 (ganeshie8):

curl=0 tells you the field is conservative so a potential function exists

ganeshie8 (ganeshie8):

go and find it

ganeshie8 (ganeshie8):

\[\vec{G} = 5xe^{5x^2+2y^2+z^2} i + 2ye^{5x^2+2y^2+z^2}j + ze^{5x^2+2y^2+z^2}k\]

ganeshie8 (ganeshie8):

if you stare at it long enough you will see that below potential function works : \[f(x,y,z) = \frac{1}{2}e^{5x^2+2y^2+z^2}\]

ganeshie8 (ganeshie8):

take partials of that function and convince yourself that you get the components of the given vector field

OpenStudy (anonymous):

so you take the integral of f(x,y,z)

ganeshie8 (ganeshie8):

find the gradient of f(x,y,z)

ganeshie8 (ganeshie8):

find the partial derivatives

OpenStudy (anonymous):

if you take that gradient you get G

ganeshie8 (ganeshie8):

Exactly! we also say G is a gradient field

ganeshie8 (ganeshie8):

all below are equivalent : 1) conservative field 2) gradient field 3) path independent field 4) potential function exists for a field 5) curl is 0

ganeshie8 (ganeshie8):

if one of them in the list is true, then eveything else in that list is also true

OpenStudy (anonymous):

so by taking the partial derivatives of f(x,y,z) you get the answer?

ganeshie8 (ganeshie8):

taking partial derivatives convinces you that below is the potential function of given vector field \[f(x,y,z) = \frac{1}{2}e^{5x^2+2y^2+z^2}\]

ganeshie8 (ganeshie8):

so the work done when you go from (3, 2, -2) to (6, 1, 7) equals : \[f(6,1,7) - f(3,2,-2)\]

ganeshie8 (ganeshie8):

evaluate ^

OpenStudy (anonymous):

so technically all you need is to find the potential function if the curl is 0

OpenStudy (anonymous):

great

ganeshie8 (ganeshie8):

Yep because setting up line integral is more painful

ganeshie8 (ganeshie8):

finding the potential function is easy

ganeshie8 (ganeshie8):

and you can find potential function ONLY when the field is conservative

ganeshie8 (ganeshie8):

potential function wont exist if the field is not conservative

OpenStudy (anonymous):

so what about part a? is using the FTFLT a good idea?

ganeshie8 (ganeshie8):

whats FTFLT ?

OpenStudy (anonymous):

fundamental theorem of line integrals

ganeshie8 (ganeshie8):

that works only for conservative fields right ?

OpenStudy (anonymous):

smooth curves

ganeshie8 (ganeshie8):

we have used fundamental theorem of line integral for part b

ganeshie8 (ganeshie8):

but it wont apply for part a because the field is not conservative

ganeshie8 (ganeshie8):

find the curl and see if you get 0 or not

OpenStudy (anonymous):

curl is (0,0,-5)

ganeshie8 (ganeshie8):

its non zero

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

that tells us that the vector field in part a is not conservative

ganeshie8 (ganeshie8):

so you cant find a potential function

ganeshie8 (ganeshie8):

that means fundamental theorem of calc for line integrals is not useful here

ganeshie8 (ganeshie8):

is that clear ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what could we use?

ganeshie8 (ganeshie8):

parameterize the path as usual

ganeshie8 (ganeshie8):

try 11.5

OpenStudy (anonymous):

well i dont have any more trys

OpenStudy (anonymous):

so for the parametric equation you get x=3+3t, y=2-t, z-2+9t

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how did you get 11.5? @ganeshie8

ganeshie8 (ganeshie8):

evaluate the line integral

ganeshie8 (ganeshie8):

x=3+3t, y=2-t, z = -2+9t dx = 3 dt dy = -1 dt dz = 9 dt

ganeshie8 (ganeshie8):

\[\int_C \vec{F}\cdot d\vec{r} = \int_C 5ydx + 2dy - dz \]

ganeshie8 (ganeshie8):

\[= \int\limits_0^1 5(2-t)(3 dt) + 2(-dt) - 9dt\]

ganeshie8 (ganeshie8):

\[= \int\limits_0^1 (5(2-t)(3 ) + 2(-1) - 9)~dt\]

ganeshie8 (ganeshie8):

simplify and evaluate the integral ^

OpenStudy (anonymous):

yeah i got this but i have a equation how do you know the limits of integration?

ganeshie8 (ganeshie8):

x=3+3t, y=2-t, z = -2+9t you are starting at (3,2,-2), so 3+3t = 3 2-t = 2 -2+9t = -2 solve t

ganeshie8 (ganeshie8):

that gives you lower bound of t

ganeshie8 (ganeshie8):

you can find the upper bound similarly

OpenStudy (anonymous):

oh ok I understand it now thank you @ganeshie8

ganeshie8 (ganeshie8):

when you parameterize a line segment like this by using the direction vector the bounds will be always (0,1) il let you figure out why..